# Learning efficient erasure protocols for an underdamped memory

### v1

@@ Abstract
Introduction — To erase one bit of information infinitely slowly requires the dissipation of $k_{\rm B}T\ln 2$ units of heat, where $k_{\rm B}$ is Boltzmann’s constant and $T$ the equilibrium temperature, a limit described by Landauer in 1961 [1, 2]. Performing erasure at finite rate incurs additional heat costs that depend on how the bit is realized and which time-dependent protocol is used [3, 4, 5, 6, 7, 8]. Experiments on a mechanical cantilever and phenomenological modeling indicate that a one-bit memory realized by an underdamped oscillator can be used to perform erasure more reliably and at higher speeds than can overdamped systems. However, two phenomena negatively affect the operation of underdamped memories: 1) residual damping losses are proportional to the velocity, and 2) the system heats up at low dissipation when heat transfer to the thermostat cannot compensate for the work done on the system. The latter energy overhead is described by an extended Landauer bound, where the temperature to consider is a weighted average of $T$ during the transformation [9, 10, 11].
@@
Harnessing the potential of underdamped systems for logic operations [12] requires finding optimal ways to perform erasure. Optimal protocols are not known because the problem is nonlinear [13, 14] – memories are bistable systems – and we want to work in the limit of fast processes, where long-time approximations [15, 16, 17] break down. In this Letter we show that evolutionary reinforcement learning [18, 19, 20, 21], a form of machine learning, can be used to find protocols for memory erasure in underdamped systems that are considerably more efficient than our best hand-designed protocols. To do so we express a time-dependent erasure protocol in the form of a deep neural network. Such an ansatz does not presuppose a functional form for the protocol, and is flexible enough to represent many different forms, including jumps and high-frequency features that appear in optimal protocols in other underdamped systems [22]. Using a simulation model (or “digital twin”) of the experiment, we apply a genetic algorithm to train the neural network in order to achieve a desired objective, which in this work is to render repeated erasures as reliable as possible, with a low energy of the memory after a single erasure as an intermediate objective. We then take the trained neural-network protocols and apply them to experiment, where we find that they behave as they do in simulation and represent a considerable improvement over our best hand-designed protocols, which were introduced in Refs. [6, 9, 10].
@@
In what follows we introduce the experimental realization of the cantilever and our simulation model of the experiment. We show that the model is an accurate representation of experiment. We then use evolutionary reinforcement learning within simulations in order to derive efficient erasure protocols, and we apply these protocols to experiment, where we observe results consistent with simulation. The trained neural-network protocols allow reliable high-speed erasures in a physical model of a one-bit memory, potentially important for next-generation computing applications. A striking benefit of those protocols is that they are also low-energy: they keep the memory cool, increasing its reliability. More generally, the combination of methods used here opens the door to the rational design of efficient protocols for a variety of physics applications.
@@
Figure 1: (a) Experimental setup: a conductive cantilever in a feedback-modulated electric field experiences (b) a tunable double-well potential [24, 25]. Eq. (1) is shown as a plain line, with experimental measurements shown as dots with error bars of one standard deviation. Logical states 0 and 1 are assigned to values $x<0$ and $x>0$, respectively. Our aim is to change the tunable parameters $x_{0},x_{1}$ as a function of time in order to render multiple erasures (resets to 0 or 1) as reliable as possible. (c) Basic erasure protocol for a reset to state 0 (top). This protocol results in the probability distribution $\rho(x,t)$ of cantilever positions in experiment (bottom). (d) Time-ordered snapshots with the potential $U(x)$, associated Boltzmann equilibrium distributions $\rho_{0}(x)$, and the distribution of cantilever positions $\rho(x)$ in experiment (green) and simulation (black dashed) [26]. The full movie of the time evolution of $U(x,t)$ and $\rho(x,t)$ corresponding to those snapshots is available in Ref. [27]. Experiment and simulation are consistent. This basic protocol is only partially reliable: about $7\%$ of trajectories end in the wrong well. In what follows we apply machine learning to the simulation in order to learn more efficient erasure protocols $[x_{0}(t),x_{1}(t)]$, and we deploy these protocols in experiment.
Using a fast feedback loop [24, 25] we can modulate the electrostatic force acting on the cantilever so that it experiences a energy $U$ parameterized by two scalars, $x_{0}$ and $x_{1}$:
$\begin{split}U(x)=&\frac{1}{2}\big{(}x-S(x-x_{0})x_{1}\big{)}^{2}\\ &+x_{0}x_{1}\big{(}S(x-x_{0})+S(x_{0})\big{)}.\end{split}$ (1)
Here $S$ is the sign function: $S(x)=-1$ if $x<0$ and $S(x)=1$ otherwise. $U(x)$ has in general a double-well form, shown in Fig. 1(b), where $x_{0}$ tunes the asymmetry and $x_{1}$ the barrier height. We start and end with a symmetric double well, and associate cantilever positions $x<0$ and $x>0$ with logical states 0 and 1, respectively. Starting in thermal equilibrium, with the cantilever in either well, we can impose a time-dependent protocol $[x_{0}(t),x_{1}(t)]$ in order to perform erasure, i.e. to bring the cantilever tip from its starting well to a specified final well in time $t_{\rm f}=2t_{0}$. We allow the protocol to act over one fundamental oscillation period $t_{0}$, and then wait in the final double-well for an additional time $t_{0}$ in order to assess the stability of the system. The time $t_{\rm f}$ is very short, smaller than the relaxation time of the system $Qt_{0}/\pi$, and optimized protocols are required in order to minimize the kinetic energy of the cantilever post-erasure and so render erasure reliable. If the post-erasure kinetic energy is too large then the cantilever will escape from the intended well, rendering erasure unreliable [11]. In what follows our aim is to determine a protocol that makes multiple erasures, each of time $t_{\rm f}$ and with no waiting time between them, as reliable as possible.
Our model of the underdamped oscillator consists of the dimensionless Langevin equation (Appendix S1)
$\ddot{x}+Q^{-1}\dot{x}=-U^{\prime}(x)+\sqrt{2Q^{-1}}\,\xi.$ (2)
Here dots and primes denote differentiation with respect to time $t$ and position $x$, respectively, and $\xi$ is a delta correlated Gaussian white noise of unit variance: $\langle\xi(t)\rangle=0$, $\langle\xi(t)^{2}\rangle=1$, and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$.
Introducing the cantilever velocity $v=\dot{x}$ and integrating (2) over the short time interval $\Delta t$ yields the update equations (Appendix S1)
@@
$v(t+\Delta t)$ $=\alpha v(t)-\left(1-\alpha\right)QU^{\prime}[x(t)]+\sqrt{1-\alpha^{2}}\xi(t),$
where $\alpha\equiv{\rm e}^{-\Delta t/Q}$. We numerically integrate Eqs. (3) from time $t=0$ to $t=t_{\rm f}$ using a timestep $\Delta t=1.09\times 10^{-4}t_{0}$, starting in equilibrium with the potential in its symmetric double-well form.
Hand-designed erasure protocols and validation of the simulation model — In Fig. 1(c) we present our basic erasure protocol [6]. As shown in Fig. 1(c-d), this merges the two wells symmetrically, translates the resulting single well to the left or right (for erasure to state 0 or 1, respectively), and then reconstitutes the double-well potential. The time-ordered snapshots in panel (d) show the potential (black), the associated Boltzmann equilibrium distributions (grey dashed), and the instantaneous distribution of cantilever positions in experiment (green) and simulation (black dashed) [26]. Our simulation results are consistent with our experiments. The cantilever positions are in general far from thermal equilibrium. From inspection of the panels at times $t/t_{\rm f}=0.7$ and 1, and from panel (c), it is clear that in several trajectories the cantilever has so much kinetic energy that it escapes from the desired well prior to the end of the quiescent period. As a result, the basic protocol achieves erasure with only about 93% success rate (in both experiment and simulation).
@@
Figure 3: Erasure protocols for a single reset to state 0 (top panels), and the resulting distributions $\rho(x,t)$ of cantilever positions in experiment. The format is the same as Fig. 1(c). We show (a) the enhanced protocol, (b) the protocol trained to maximize the success rate and minimize the total energy upon performing a single erasure, and (c) the protocol trained to maximize the success rate of 100 consecutive erasures. In Fig. S7 in the appendix we show that distributions $\rho(x,t)$ from simulation and experiment are consistent. Movies of the time evolution of $U(x,t)$ and $\rho(x,t)$ for all protocols are available in Ref. 27.
@@
In Fig. 2(a) we show that the basic protocol fails completely when applied to 100 consecutive erasures (with the target 0 or 1 chosen randomly with equal likelihood): the probability $P_{\rm s}(100)$ of observing 100 consecutive successful erasures is zero, in both simulation and experiment [28]. The excess kinetic energy possessed by the cantilever following each erasure quickly builds and renders erasures completely unreliable. In Fig. 2(b) we illustrate the work distribution associated with the basic erasure protocol in simulation and experiment, confirming again the accuracy of the simulation model. This distribution is very wide, with a mean value $\langle W\rangle=15.9$ far above the Landauer bound. With such a short $t_{\rm f}$, the basic protocol is both unreliable and expensive.
@@
In Fig. 3(a) we show our best hand-designed protocol, an enhanced version of the basic protocol with delta function-like kicks at four points along the trajectory [10]. These kicks are motivated by the finding that optimal translation of an underdamped particle in an harmonic potential requires delta function impulses that abruptly change the velocity of the particle [22]. In our case the enhanced protocol outperforms the basic protocol to a considerable degree, achieving a probability of success for one erasure of $P_{\rm s}(1)=99.2\%$. However, for 100 consecutive erasures the success rate drops sharply, to $P_{\rm s}(100)=13\%$ [Fig. 2(a)]. The work distribution is much narrower than that of the basic protocol, with the average work decaying to $\langle W\rangle=6.5$ [Fig. 2(b)]. Part of the overhead to Landauer’s bound originates from the kicks around $t/t_{\rm f}=0.25$. These are sometimes in the wrong direction, creating the side lobes around the ideal trajectory in Fig. 3(a). Some of these side-lobe trajectories eventually escape to the wrong well.
@@
Learning optimized protocols — We can use the simulation model to produce an improved erasure protocol. Following Ref. 29, we encode the time-dependent protocol $[x_{0}(t),x_{1}(t)]$ in the form of a deep neural network, and train it using a genetic algorithm [18, 30] to achieve a desired goal (Appendix S3). This process is a form of evolutionary reinforcement learning. We have carefully benchmarked this procedure using overdamped passive-matter [29, 31] and active-matter systems [32], and in Appendix S2 we show that it performs well in the case of an underdamped system whose optimal translation protocol is known exactly [22].
@@
In Fig. 3(c) we show the protocol produced by retraining the protocol of panel (b) to maximize the order parameter $\phi=P_{\rm s}(100)$, the success rate for 100 consecutive erasures (averaged over $10^{4}$ independent realizations of the process). It is subtly different from the previous protocol, but performs considerably better on consecutive erasures: we measure success rates for 100 consecutive erasures of 90.9 % (simulation) and 92.1 % (experiment). That is, the learned protocol is essentially as effective when applied to 100 consecutive erasures as the basic protocol is for a single erasure.
Mechanism underpinning the success of the learned protools — The enhanced and learned protocols achieve a single erasure with high probability, 99.2% and 99.9%, respectively, and the distribution of cantilever positions at time $t_{\rm f}$ is similar for the two protocols (see Fig. 3 and Fig. S8 in the appendix). The main difference between the two protocols is the kinetic energy of the cantilever after erasure. In Fig. 4(a) we show distributions of cantilever velocities after one erasure for the basic protocol, enhanced protocol, and learned protocol of Fig. 3(b), superposed with the equilibrium distribution at temperature $T$. The basic protocol has a long tail of high-energy values. The enhanced protocol is better, but the learned protocol is better still: the evolutionary optimization process was instructed to maximize erasure probability and minimize total energy (kinetic plus potential), and it has done this successfully.
@@
Conclusions — We have used evolutionary reinforcement learning applied to a simulation model to identify efficient erasure protocols for an underdamped mechanical cantilever. We have shown that the same protocols when used in experiment are considerably more efficient than our best hand-designed protocols. In this case the simulation model is an accurate representation of our experiment. If it were not, there exists the possibility of applying the learning algorithm directly to experiment, because all the information needed to train the neural network is accessible in experiment. This combination of methods therefore opens the door to the rational design of efficient protocols for applied physics generally.
The data supporting this study is available in Ref. 33. Code for doing evolutionary optimization of cantilever protocols can be found at Ref. 34.
Acknowledgments — This work has been partially funded by project ANR-22-CE42-0022. SW performed work as part of a user project at the Molecular Foundry at Lawrence Berkeley National Laboratory, supported by the Office of Basic Energy Sciences of the U.S. Department of Energy under Contract No. DE-AC02–05CH11231. SW was partially supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences Data, Artificial Intelligence and Machine Learning at DOE Scientific User Facilities program under Award Number 34532 (a digital twin for in silico spatiotemporally-resolved experiments).

### v2

@@ I Introduction
To erase one bit of information infinitely slowly requires the dissipation of $k_{\rm B}T\ln 2$ of heat, where $k_{\rm B}$ is Boltzmann’s constant and $T$ the equilibrium temperature, a limit described by Landauer in 1961 [1, 2]. Performing erasure at finite rate incurs additional heat costs that depend on how the bit is realized and which time-dependent protocol is used [3, 4, 5, 6, 7, 8]. Experiments on a mechanical cantilever and phenomenological modeling indicate that a one-bit memory realized by an underdamped oscillator can be used to perform erasure more reliably and at higher speeds than can overdamped systems. However, two phenomena negatively affect the operation of underdamped memories: 1) residual damping losses are proportional to the velocity, and 2) the system heats up at low dissipation when heat transfer to the thermostat cannot compensate for the work done on the system. The latter energy overhead is described by an extended Landauer bound, where the temperature to consider is a weighted average of $T$ during the transformation [9, 10, 11].
@@
Harnessing the potential of underdamped systems for logic operations [12] requires finding optimal ways to perform erasure. Optimal protocols are not known because the problem is nonlinear [13, 14] – memories are bistable systems – and we want to work in the limit of fast processes, where long-time approximations [15, 16, 17] break down. In this article we show that evolutionary reinforcement learning [18, 19, 20, 21], a form of machine learning, can be used to find protocols for memory erasure in underdamped systems that are considerably more efficient than our best hand-designed protocols. To do so we express a time-dependent erasure protocol in the form of a deep neural network. Such an ansatz does not presuppose a functional form for the protocol, and is flexible enough to represent many different forms, including jumps and high-frequency features that appear in optimal protocols in other underdamped systems [22]. Using a simulation model (or “digital twin”) of the experiment, we apply a genetic algorithm to train the neural network in order to achieve a desired objective, which in this work is to render repeated erasures as reliable as possible, with a low energy of the memory after a single erasure as an intermediate objective. We then take the trained neural-network protocols and apply them to experiment, where we find that they behave as they do in simulation and represent a considerable improvement over our best hand-designed protocols, which were introduced in Refs. [6, 9, 10].
@@
In what follows we introduce the experimental realization of a 1-bit memory with a cantilever and our simulation model of the experiment. We show that the model is an accurate representation of experiment. We then use evolutionary reinforcement learning within simulations in order to derive efficient erasure protocols, and we apply these protocols to experiment, where we observe results consistent with simulation. The trained neural-network protocols allow reliable high-speed erasures in a physical model of a one-bit memory, potentially important for next-generation computing applications. A striking benefit of those protocols is that they are also low-energy: they keep the memory cool, increasing its reliability. The combination of methods used here and the ability to deploy protocols in experiments opens the door to the rational design of hardware and software for deterministic $k_{\rm B}T$-scale computing, and more generally to efficient protocols for a variety of physics applications.
II Experimental apparatus and simulation model
Figure 1: (a) Experimental setup: a conductive cantilever experiences an electrostatic force $F_{\mathrm{el}}$ due to the voltage difference $V-V_{0}$ with a facing electrode. The deflection $x$ of the cantilever, measured by interferometry, is compared to a threshold $x_{0}$ by the feedback loop setting $V=\pm V_{1}$. (b) This results in a tunable double-well potential $U$ [23, 24], that can be measured via the Boltzmann distribution $\rho_{0}(x)$ of the cantilever deflection in equilibrium: $U(x)=-\ln\rho_{0}(x)$. We plot this measured potential as dots, with error bars of one standard deviation, on top of the line denoting Eq. (1). Logical states 0 and 1 are assigned to values $x<0$ and $x>0$, respectively. Our aim is to change the tunable parameters $x_{0},x_{1}$ as a function of time in order to render multiple erasures (resets to 0 or 1) as reliable as possible. (c) Basic erasure protocol for a reset to state 0 (top). This protocol results in the distribution $\rho(x,t)$ in experiment (bottom). (d) Time-ordered snapshots with the potential $U(x,t)$, associated Boltzmann equilibrium distributions $\rho_{0}(x)$, and the distribution of cantilever positions $\rho(x,t)$ in experiment (green) and simulation (black dashed) [25]. The full movie of the time evolution of $U(x,t)$ and $\rho(x,t)$ corresponding to those snapshots is available as an ancillary file [26]. Experiment and simulation are consistent. This basic protocol is only partially reliable: about $7\%$ of trajectories end in the wrong well. In what follows we apply machine learning to the simulation in order to learn more efficient erasure protocols $[x_{0}(t),x_{1}(t)]$, and we deploy these protocols in experiment.
@@ II Experimental apparatus and simulation model
Using a fast feedback loop [23, 24] we can modulate the electrostatic force acting on the cantilever so that it experiences a virtual energy potential $U(x,t)$ that is parameterized by two scalars, $x_{0}(t)$ and $x_{1}(t)$:
$\begin{split}U(x,t)=&\frac{1}{2}\big{(}x-S[x-x_{0}(t)]x_{1}(t)\big{)}^{2}\\ &+x_{0}(t)x_{1}(t)\big{[}S(x-x_{0}(t)]+S[x_{0}(t)]\big{)}.\end{split}$ (1)
Here $S$ is the sign function, where $S[x]=-1$ if $x<0$, and $S[x]=1$ otherwise. $U(x,t)$ has in general a double-well form, shown in Fig. 1(b), where $x_{0}$ tunes the asymmetry and $x_{1}$ the barrier height. We start and end with a symmetric double well, and associate cantilever positions $x<0$ and $x>0$ with logical states 0 and 1, respectively. Starting in thermal equilibrium, with the cantilever in either well, we can impose a time-dependent protocol $[x_{0}(t),x_{1}(t)]$ in order to perform erasure, i.e. to bring the cantilever tip from its starting well to a specified final well in time $t_{\rm f}=2t_{0}$. We allow the protocol to act over one fundamental oscillation period $t_{0}$, and then wait in the final double-well for an additional time $t_{0}$ in order to assess the stability of the system. The time $t_{\rm f}$ is very short, smaller than the relaxation time of the system $Qt_{0}/\pi$, and optimized protocols are required in order to minimize the kinetic energy of the cantilever post-erasure and so render erasure reliable. If the post-erasure kinetic energy is too large then the cantilever will escape from the intended well, rendering erasure unreliable [11]. In what follows our aim is to determine a protocol that makes multiple erasures, each of time $t_{\rm f}$ and with no waiting time between them, as reliable as possible.
Our model of the underdamped oscillator consists of the dimensionless Langevin equation (see derivation in Appendix A)
$\ddot{x}+Q^{-1}\dot{x}=-U^{\prime}(x,t)+\sqrt{2Q^{-1}}\,\xi.$ (2)
Here dots and primes denote differentiation with respect to time $t$ and position $x$, respectively, and $\xi$ is a Gaussian white noise satisfying $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$.
Introducing the cantilever velocity $v=\dot{x}$ and integrating Eq. (2) over the short time interval $\Delta t$ yields the update equations (details in Appendix A)
@@
$v(t+\Delta t)$ $=\alpha v(t)-\left(1-\alpha\right)QU^{\prime}[x(t)]+\sqrt{1-\alpha^{2}}\tilde{\xi}(t),$
where $\alpha\equiv{\rm e}^{-\Delta t/Q}$, and $\tilde{\xi}(t)$ is randomly drawn from a centered normal distribution of unit variance. We numerically integrate Eqs. (3) from time $t=0$ to $t=t_{\rm f}$ using a timestep $\Delta t=1.09\times 10^{-4}t_{0}$, starting in equilibrium with the potential in its symmetric double-well form.
III Hand-designed erasure protocols and validation of the simulation model
In Fig. 1(c) we present our basic erasure protocol [6]. As shown in Fig. 1(c-d), this merges the two wells symmetrically, translates the resulting single well to the left or right (for erasure to state 0 or 1, respectively), and then reconstitutes the double-well potential. The time-ordered snapshots in panel (d) show the potential (black), the associated Boltzmann equilibrium distributions (grey dashed), and the instantaneous distribution of cantilever positions in experiment (green) and simulation (black dashed) [25]. Our simulation results are consistent with our experiments. The cantilever positions are in general far from thermal equilibrium. From inspection of the panels at times $t/t_{\rm f}=0.7$ and 1, and from panel (c), it is clear that in several trajectories the cantilever has so much kinetic energy that it escapes from the desired well prior to the end of the quiescent period. As a result, the basic protocol achieves erasure with only about 93% success rate (in both experiment and simulation).
@@ III Hand-designed erasure protocols and validation of the simulation model
Figure 3: Erasure protocols for a single reset to state 0 (top panels), and the resulting distributions $\rho(x,t)$ of cantilever positions in experiment. The format is the same as Fig. 1(c). We show (a) the enhanced protocol, (b) the protocol trained to maximize the success rate and minimize the total energy upon performing a single erasure, and (c) the protocol trained to maximize the success rate of 100 consecutive erasures. In Fig. 7 in the appendix we show that distributions $\rho(x,t)$ from simulation and experiment are consistent. Movies of the time evolution of $U(x,t)$ and $\rho(x,t)$ for all protocols are available as ancillary files [26].
@@
In Fig. 2(a) we show that the basic protocol fails completely when applied to 100 consecutive erasures (with the target 0 or 1 chosen randomly with equal likelihood): the probability $P_{\rm s}(100)$ of observing 100 consecutive successful erasures is zero, in both simulation and experiment [28]. The excess kinetic energy possessed by the cantilever following each erasure quickly builds and renders erasures completely unreliable. For all trajectories, we compute the work $W$ performed on the system during the erasure using the usual stochastic thermodynamics convention [29], $W=\int\partial_{t}U(x,t)dt$. In the experiment, the feedback loop has no energetic contribution, because its only role is to construct the virtual potential [23, 24]. In Fig. 2(b) we illustrate the work distribution associated with the basic erasure protocol in simulation and experiment, confirming again the accuracy of the simulation model, and underlining the fact that feedback used to create the virtual potential – which is not considered in the simulation model – has no energetic impact upon erasure. The work distribution is very broad, with a mean value $\langle W\rangle=15.9$, far above the Landauer bound. With such a short $t_{\rm f}$, the basic protocol is both unreliable and expensive.
@@
In Fig. 3(a) we show our best hand-designed protocol, an enhanced version of the basic protocol with delta function-like kicks at four points along the trajectory [10]. These kicks are motivated by the finding that optimal translation of an underdamped particle in a harmonic potential requires delta function impulses that abruptly change the velocity of the particle [22]. In our case the enhanced protocol outperforms the basic protocol to a considerable degree, achieving a probability of success for one erasure of $P_{\rm s}(1)=99.2\%$. However, for 100 consecutive erasures the success rate drops sharply, to $P_{\rm s}(100)=13\%$ [Fig. 2(a)]. The work distribution is much narrower than that of the basic protocol, with the average work decaying to $\langle W\rangle=6.5$ [Fig. 2(b)]. Part of the overhead to Landauer’s bound originates from the kicks around $t/t_{\rm f}=0.25$. These are sometimes in the wrong direction, creating the side lobes around the ideal trajectory in Fig. 3(a). Some of these side-lobe trajectories eventually escape to the wrong well.
IV Learning optimized protocols
@@
We can use the simulation model to produce an improved erasure protocol. Following Ref. 30, we encode the time-dependent protocol $[x_{0}(t),x_{1}(t)]$ in the form of a deep neural network, and train it using a genetic algorithm [18, 31] to achieve a desired goal (more details in Appendix C). This process is a form of evolutionary reinforcement learning. We have carefully benchmarked this procedure using overdamped passive-matter [30, 32] and active-matter systems [33], and in Appendix B we show that it performs well in the case of an underdamped system whose optimal translation protocol is known exactly [22].
In Fig. 3(c) we show the protocol produced by retraining the protocol of panel (b) to maximize the order parameter $\phi=P_{\rm s}(100)$, the success rate for 100 consecutive erasures (averaged over $10^{4}$ independent realizations of the stochastic process to ensure low statistical uncertainty). It is subtly different from the previous protocol, but performs considerably better on consecutive erasures: we measure success rates for 100 consecutive erasures of 90.9 % (simulation) and 92.1 % (experiment). That is, the learned protocol is essentially as effective when applied to 100 consecutive erasures as the basic protocol is for a single erasure.
The excellent agreement between the experiments and the numerical simulations on its digital twin show that the learned protocols are robust to non-idealities of the system, such as the description of the mechanical system and its frequency-dependent damping [34, 35] by a simple harmonic resonator. Using the dimensionless equations for the dynamics, we can in particular simply rescale the protocols for any variation of the resonant frequency $f_{0}$ or of the stiffness $k$. Small variations (of the order of $10\%$) of the quality factor $Q$ or final time $t_{\rm f}$ have no noticeable effect on the learned protocols, but larger ones would require a retraining.
V Mechanism underpinning the success of the learned protocols
The enhanced and learned protocols achieve a single erasure with high probability, 99.2% and 99.9%, respectively, and the distribution of cantilever positions at time $t_{\rm f}$ is similar for the two protocols (see Fig. 3 and Fig. 8 in Appendix D). The main difference between the two protocols is the kinetic energy of the cantilever after erasure. In Fig. 4(a) we show distributions of cantilever velocities after one erasure for the basic protocol, enhanced protocol, and learned protocol of Fig. 3(b), superposed with the equilibrium distribution at temperature $T$. The basic protocol has a long tail of high-energy values. The enhanced protocol is better, but the learned protocol is better still: the evolutionary optimization process was instructed to maximize erasure probability and minimize total energy (kinetic plus potential), and it has done this successfully.
@@ V Mechanism underpinning the success of the learned protocols
As illustrated in Fig. 4(b), the memory reaches a non equilibrium steady state after only a few cycles. The choice of 100 consecutive erasures is thus more than sufficient to ensure that the resulting protocol can then be applied to a large number of consecutive erasures. The corresponding improvement in success rate over $N$ consecutive erasures done using the different protocols can be determined by extrapolating each line in Fig. 2(a), assuming an exponential decay of $P_{\rm s}$ with $N$.
VI Conclusions
We have used evolutionary reinforcement learning applied to a simulation model to identify efficient erasure protocols for an underdamped mechanical cantilever. We have shown that the same protocols when used in experiment are considerably more efficient than our best hand-designed protocols, essentially solving the heating problem that plagues repeated fast erasures. In this case the simulation model is an accurate representation of our experiment. If it were not, there exists the possibility of applying the learning algorithm directly to experiment, because all the information needed to train the neural network is accessible in experiment. This combination of methods therefore opens the door to the rational design of $k_{\rm B}T$-scale hardware and their operating algorithms, and more generally to efficient protocols for applied physics.
The data supporting this study is available in Ref. 36. Code for doing evolutionary optimization of cantilever protocols can be found at Ref. 37.
Acknowledgements.
This work has been partially funded by project ANR-22-CE42-0022. SW performed work as part of a user project at the Molecular Foundry at Lawrence Berkeley National Laboratory, supported by the Office of Basic Energy Sciences of the U.S. Department of Energy under Contract No. DE-AC02–05CH11231. SW was partially supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences Data, Artificial Intelligence and Machine Learning at DOE Scientific User Facilities program under Award Number 34532 (a digital twin for in silico spatiotemporally-resolved experiments).